# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ Abstract
Data-driven cyberthreat detection has become a crucial defense technique in modern cybersecurity. Network defense, supported by Network Intrusion Detection Systems (NIDSs), has also increasingly adopted data-driven approaches, leading to greater reliance on data. Despite its importance, data scarcity has long been recognized as a major obstacle in NIDS research. In response, the community has published many new datasets recently. However, many of them remain largely unknown and unanalyzed, leaving researchers uncertain about their suitability for specific use cases.
@@ 1 Introduction
An efficient IDS maximizes attack detection (ideally 100%) while minimizing the false alarm rate, i.e., legitimate samples flagged as malicious (ideally 0%). The performance of detection systems thus needs to be evaluated, analyzed, and compared using datasets containing malicious and legitimate samples. For this reason, the community relies on public intrusion datasets to research and develop new IDSs. The importance of data has been further emphasized by the rise of data-driven approaches for intrusion detection, most notably machine learning (ML) [8].
@@
    We perform a systematic, up-to-date comparative survey of public datasets for NIDS research and development published until 2023 (inclusive). This will introduce readers to the newest trends and possibilities for conducting experiments and method benchmarks. To the best of our knowledge, this survey covers the largest number of publicly available NID datasets (89).
@@
    In order to aid in data selection, we extract 13 properties for each dataset by manually downloading and exploratory analyzing the data. We also study the popularity of datasets in state-of-the-art NIDS research. To enhance accessibility, we share our analysis notebooks on GitHub [^1] https://github.com/xGoldy/nid-datasets, allowing others to review the data structure without downloading it.
@@
    Finally, we advocate the correct creation and usage of the data by providing data handling recommendations.
The rest of the paper is composed as follows: We define the survey’s scope and methodology in Section 2. Section 3 elaborates on the NID domain-specific properties and their impact on data. Section 4 describes 13 collected properties used to compare the datasets. The main part of this paper – the data survey, is presented in Section 5, listing datasets in a tabular format, categorizing them, and briefly elaborating on their properties and potential use cases. This section also analyzes dataset popularity and trending developments in NID data research. We provide recommendations for data selection, creation, and usage in Section 6. Finally, sections 7 – 9 present related work, conclude the paper, and outline future research directions.
@@ 2 Survey Methodology
In order to provide comprehensive, unbiased, and reproducible results, we used the Systematic Literature Review (SLR) for computer science [15] methodology to conduct the survey. This section briefly describes the survey’s scope and research questions, whereas details are discussed in A.
@@ 2.1 Scope of the Survey
Our paper aims to provide a holistic view of the network intrusion datasets and their characteristics. Therefore, we do not narrow down on any specific environment or intrusion type but include datasets containing network traffic from various environments like the Internet of Things (IoT), Cyber-Physical Systems (CPSs), Software Defined Networks (SDNs), or Critical infrastructure Supervisory Control and Data Acquisition (SCADA).
Although not restricted to a particular environment, we strictly include only datasets containing network traces. Therefore, datasets without activity observable on a computer network (i.e., packets or flows) were excluded. As a result, the survey did not consider datasets that included only environment-specific features like sensor values typical for CPS or SCADA.
Given the mentioned scope, we emphasize that this survey is not supposed to be an exhaustive list of all available IDS datasets and, thus, not a sole data reference point for specific use cases. Instead, we encourage looking for other surveys focused on specific cybersecurity subdomains.
@@ 3 Domain-Specific Properties and Their Impact on Data
Despite its rapid adoption in research, ML penetrated into practical cybersecurity solutions much slower than in other domains, such as computer vision or natural language processing, which have been actively using ML for over a decade [16]. In fact, only a fraction of real-world network monitoring and threat detection teams rely on machine learning nowadays [17].
The discrepancy between academic research and the actual deployments of ML-based NIDSs was discussed by Sommer and Paxson [12], who noticed that operators were reluctant to adopt ML-based anomaly detectors despite high research efforts within the area. This fact was attributed to fundamental differences between NID and other ML-application domains, driven by factors like very high cost of errors, enormous data variability, semantic gap (ML explainability issues), and difficulties with evaluation – including the lack of benchmark data. Although published more than 10 years ago, most of these problems arose from the specific properties of cybersecurity and computer networks domains, affecting the research and development efforts to this day. This section elaborates on these issues and their impact on overall NID data availability and quality.
@@ 3.1 When Hell Comes to Earth: Unpleasant Properties of the Network Intrusion Detection Domain
    Network traffic diversity and dynamicity (intra-network variability): Computer networks are highly dynamic by nature. Network environments constantly change as new devices are connected to the network, new applications and protocols are developed, or even new networking paradigms, such as the Internet of Things – IoT, are introduced. The temporal aspect further introduces periodicity patterns based on day-and-night cycles and office hours [19].
@@
    Adversarial environment: In addition to natural networks’ dynamicity, the NID domain has to assume the implicit presence of adversaries. Due to the endless cybersecurity arms race, the attacks continuously evolve, and new vulnerabilities are found. Nevertheless, the attackers might also attempt to bypass detection via adversarial samples crafted with obfuscation techniques or tiny perturbations of the input data [20].
@@
    In NID, even domain experts might struggle to verify ground truth, whereas attack detectors might return different and even contradicting results [25]. These facts also disable crowdsourcing annotation, a common way to obtain large corpora of labeled data [26]. Finally, even all the traffic of infected hosts cannot be considered malicious, as some might correspond to legitimate activities (e.g., ARP messages) [22]. For these reasons, accurately labeling real-world data is difficult and often infeasible.
@@
    Data confidentiality: Captures from real networks might contain highly sensitive and confidential information, such as IP addresses or packet payloads, significantly hampering their sharing. Although anonymization techniques [27] like IP address remapping or payload stripping can be used, they decrease the data’s value – e.g., payload analyzers NIDSs cannot be used.
Considering the labeling and data sharing issues, the lack of quality real-world data becomes understandable. However, the problem is not solved even when such data are collected and labeled. Since data is captured in a particular time in the past, future changes in the underlying distribution (data drift) or a relationship between input features and the output labels (concept drift) violate the i. i. d. assumption. This behavior, called model drifting, causes degradation of the prediction model’s performance in time. Since NID is adversarial and highly dynamic, model drifting, amplified by the low tolerance of errors, is considered a major obstacle in ML utilization in real-world scenarios. Although paradigms like incremental [28], lifelong [29], or active learning [30] utilizing the newly observed data for updating the model were proposed to tackle drifting, applying ML for real-world NIDSs still remains heavily constrained.
@@ 3.2 Limitations of Existing Datasets
    Timeliness: The direct consequence of constantly evolving network and threat landscapes is data timeliness. Generally, it cannot be assumed that the behavior collected during a certain period will remain immutable over time [31]. Therefore, static datasets become less representative of real-world scenarios as they age, affecting the soundness of the evaluation results. Although continual updates with the newest traffic patterns would mitigate this issue, immutable (static) datasets are required for fair benchmark comparisons between systems.
@@
    Testbed limitations: Due to the limitations of real-world data, most public datasets contain emulated traffic captured in a lab testbed environment. Despite some benefits of such an approach, its data are not guaranteed to resemble realistic network behavior and threats [34]. The lack of realism in benchmark data is considered one of major obstacles in deploying anomaly-based NIDSs in real-world production environments [12, 31]. For instance, Catillo et al. [35] have shown that many attacks from DoS public datasets do not resemble real-world characteristics due to being inefficient (not achieving their malicious intent) or assuming non-realistic environment configurations.
@@
    Class imbalance: Modern machine learning classifiers are trained to maximize predictive accuracy by minimizing misclassification errors [36]. Such a setting usually assumes roughly balanced classes, i.e., with similar prior probabilities. However, cyberthreat detection domains like NID typically deal with a huge imbalance in favor of the negative (benign) class, as only a few attack instances are observed along a vast amount of non-malicious communication. Although testing data should resemble real-world distribution to prevent spatial bias [37], training a supervised classifier should be done on approximately balanced data. Nevertheless, most datasets are imbalanced, leaving to address the issue by their users [38].
@@
    Documentation: The dataset’s accompanying metadata and documentation are crucial in determining the data value and its suitability for particular scenarios. However, datasets often lack the documentation completely or omit necessary details. For this reason, the soundness of conclusions drawn by the data users is significantly hampered.
@@
    Lack of explicit train-test splits: Many datasets are not explicitly split for training and testing. As a result, data users usually create arbitrary splits, causing their methods to be incomparable, as each may be trained and evaluated on a different data subset. Furthermore, applying methods like cross-validation or partitioning the data without respecting time dependencies might lead to temporal bias and inflated evaluation performance [14, 37].
@@
    Feature set discrepancies: As shown in our survey, most dataset authors propose their own feature sets, typically based on network flows. While some provide only 10 – 15 essential features, others might offer over 100 advanced statistics. Several newer datasets adopt the CICFlowMeter tool [39] for feature extraction. An attempt to unify feature sets was made by Sarhan et al. [40, 41], who proposed extracting 43 NetFlow v9 features using nProbe [42] flow collector. Despite these efforts, the community has not yet agreed on a unified, widely accepted feature set, hampering the transferability and generalizability between datasets with different features.
@@
    Limited scope: Due to the diversity of network technologies and cyberthreats, assembling a comprehensive dataset that covers the entire NID landscape is infeasible. While specialized datasets (e.g., IoT) have been released recently, many scenarios and intrusions remain underrepresented. In particular, Hindy et al. [44] note that public datasets cover only 33% of known attacks. Given that new attacks and vulnerabilities continually emerge, a NIDS benchmarked on a static dataset will never be able to fully comprehend all potential threats in the wild.
@@
    Methodically wrong data handling: This limitation originates primarily from human error in data collection and processing. A common issue we encountered relates to improper timestamping, as some datasets contain incorrect timestamps or time inconsistencies of events (e.g., attacks) with the documentation. We also observed disordered packets, undocumented capture gaps, and capture interruptions, resulting in improperly terminated data files. Data processing issues include missing values, duplicated features or entries, and zero-variance features (e.g., all zeros). Some datasets also suffer from imprecise or incorrect labeling, as shown by Liu et al. [45] on CIC-IDS data. These flaws are often unreported, creating ambiguity for users and undermining the soundness of their conclusions.
While many discussed limitations arise from domain-specific properties and are challenging to address, others (documentation, lack of explicit splits, feature set discrepancies, reproducibility, and data handling errors) are entirely within the control of dataset authors. Even if a limitation cannot be fully mitigated, properly documenting it reduces ambiguity and greatly enhances data usability. For this reason, we consider documentation a crucial pillar of every dataset. However, most do not discuss any limitations at all. Supposing that authors do not release inherently flawed records on purpose, we theorize that they often neglect thorough analysis of their collected data.
From the perspective of data users, understanding these limitations aids in efficient data selection, proper handling, and avoiding evaluation errors. In Section 6, we provide recommendations to minimize or entirely prevent the impact of limitations.
@@ 4 Network Intrusion Data and Its Properties
Depending on the dataset goal and capture attributes, we can characterize network intrusion data by various properties that influence their use cases and the soundness of methods utilizing them. In this survey, we collected 13 essential properties to provide an unbiased, comprehensive view and comparison of available datasets for NIDS purposes. A portion of them was adopted from Ring et al.’s survey [13].
We split the extracted information into five categories: General information, Nature of data, Data volume, Network properties, and Evaluation. This section elaborates on these categories and individual properties to aid in data understanding, selection, as well as support the evaluation of datasets not included in this survey in future research. Their summary is provided in Table 1.
Table 1: Summary of 13 collected properties of the surveyed NID datasets split into five categories. The table briefly describes each property, and the fourth column (“Possible values”) is a comma-separated list of values a particular property can attain. Curly brackets symbolize a set from which a single value can be drawn, whereas $\times$ is a Cartesian product of such sets in cases when multiple elements define a single property value. Ampersand (“&”) symbol is occasionally used to combine multiple values within a single entry.
@@ 4.1 General Information
When choosing a dataset, we are primarily interested in what scenarios are included in it, as well as other essential information like when the data was collected, what attacks are present, and what environment was used for the capture:
Year: The year(s) or period of data collection. As many datasets are published months or years after the collection, reporting the actual collection date helps to determine the captured threat landscape and traffic characteristics (e.g., popular network application). For continuously ongoing captures, e.g., MAWILab [46], we mark the value as Now.
Focus: This property summarizes the dataset scenario and its potential use cases. Based on the focus, we distinguish between general- and special-purpose datasets. General-purpose datasets (General) are captured in traditional local area or enterprise networks with various intrusive scenarios. Depending on the nature of the data, we include IDS (both host and network information) or NIDS (only network data are present) identifiers. In contrast, special-purpose datasets focus on a particular attack like DDoS or a specific environment such as IoT.
Normal: Indicates whether normal (benign) traffic is present and how it was generated. If absent, the property’s value is no. Otherwise, we distinguish between data obtained from real, emulated (emul), or synthetic (synth) environments. As various definitions of these terms are found in the literature (especially, emulated and synthetic terms are often used interchangeably), for the purpose of this survey, we define these terms as follows:
  • real – Traffic originating from real-world operational networks captured by monitoring tools like tcpdump or nfdump. It resembles genuine network interactions with purposes beyond traffic generation. Because it contains real information, it must be anonymized. Additionally, real data cannot be declared benign with absolute confidence, as undetected attacks might be present. Instead, we refer to it as background traffic without evident anomalous or malicious behavior.
@@
In contrast, emulated and synthetic traffic offer greater flexibility, is cheaper to collect, and labeling can be done with higher confidence. However, as outlined in Section 3.2, their crucial limitation is realism. Due to the complexity of network traffic, replicating real-world scenarios in a laboratory environment is very challenging. Although several works have tried to tackle this issue, e.g., $\beta$-profile system [50], realistic simulation of network behavior remains an important research direction [44].
@@
The attack origin environment is defined the same as for normal traffic – real, emul-{p, v, pv}, and synth. Real attacks correspond to traffic with actual malicious intent captured in real-world production networks, i.e., by detecting and sampling an ongoing attack, as in the Booters dataset [53]. Emulated attacks are launched in controlled environments just to simulate attacks’ behavior, usually with the same tools as attackers (e.g., in Kali Linux [54]). Finally, synthetic attacks were not captured in a network but rather generated by specialized algorithms.
@@
The environment origin is followed by a value in brackets indicating the number of distinguishable attacks. This usually refers to the number of malicious classes clearly separable in the data (either on a feature-based level based on labels or via raw packet filtering). For datasets like CUPID [56] that declare multiple attacks but provide only binary labels, we list the value as one, i.e., distinguishable anomalies. Datasets with unspecified attack counts, e.g., MAWILab [46], have no associated value.
In the third part, we collect the types of performed attacks. As existing taxonomies on network anomalies and attacks [44, 57, 58, 59, 60, 61] were unsuitable for our purposes – simple yet rigorous classification for any NID dataset, we utilize a custom scheme of common attacks in network intrusion datasets, as shown in Table 2. Following this scheme, each attack type corresponds to an acronym (e.g., “D” for a (D)DoS attack), with multiple types forming a single string.
@@
Table 2: The classification scheme used for listing attacks in the reviewed datasets. Each attack category corresponds to an acronym, such as “B" for Brute-force. Such a string of acronyms (e.g., BCO) then represents the presence of given types of attacks within the dataset.
@@
Botnet C&C traffic C Command and Control (C&C) traffic is used by infected machines (bots) to communicate and receive orders from the master server within the client-server botnet architectures.
@@
Vulnerability exploitation V Exploration of various vulnerabilities like SQL injection or buffer overflow typically aims to gain remote access or escalate privileges onto the target machine.

### v2

@@ Abstract
Data-driven cyberthreat detection has become a crucial defense technique in modern cybersecurity. Network defense, supported by Network Intrusion Detection Systems (NIDSs), has also increasingly adopted data-driven approaches, leading to greater reliance on data. Despite the importance of data, its scarcity has long been recognized as a major obstacle in NIDS research. In response, the community has published many new datasets recently. However, many of them remain largely unknown and unanalyzed, leaving researchers uncertain about their suitability for specific use cases.
@@ 1 Introduction
An efficient IDS maximizes attack detection (ideally 100%) while minimizing the false alarm rate, i.e., legitimate samples flagged as malicious (ideally 0%). The performance of detection systems thus needs to be evaluated, analyzed, and compared using datasets containing malicious and legitimate samples. For this reason, public intrusion datasets play an important role in IDS research and development [8]. The importance of data has been further emphasized by the rise of data-driven approaches for intrusion detection, most notably machine learning (ML) [9].
@@
    We conduct a systematic, up-to-date comparative survey of public datasets for NIDS research and development published until 2023 (inclusive). This will introduce readers to the latest trends and possibilities for conducting experiments and method benchmarks. To the best of our knowledge, this survey covers the largest number of publicly available NID datasets (89).
@@
    In order to aid in data selection, we extract 13 properties for each dataset by manually downloading and analyzing the data. We also study the popularity of datasets in state-of-the-art NIDS research. To enhance accessibility, we share our analysis notebooks on GitHub [^1] https://github.com/xGoldy/nid-datasets, enabling others to review the data structure without downloading it.
@@
    Finally, we advocate the correct creation and usage of the data by discussing data handling recommendations.
The rest of the paper is composed as follows: We define the survey’s scope and methodology in Section 2. Section 3 elaborates on the NID domain-specific properties and their impact on data. Section 4 describes 13 collected properties used to compare the datasets. The main part of this paper – the data survey, is presented in Section 5, listing datasets in a tabular format, categorizing them, and briefly elaborating on their properties and potential use cases. This section also analyzes dataset popularity and trending developments in NID data research. We discuss recommendations for data selection, creation, and usage in Section 6. Finally, sections 79 outline future research directions, discuss related work, and conclude the paper.
@@ 2 Survey Methodology
In order to ensure comprehensive, unbiased, and reproducible results, we followed the Systematic Literature Review (SLR) methodology for computer science [16] to conduct our survey. This section outlines the survey’s scope and research questions, while further details are provided in A.
@@ 2.1 Scope of the Survey
Our paper aims to provide a holistic view of the network intrusion datasets and their properties. Therefore, we do not narrow down on any specific environment or intrusion type but include datasets containing network traffic from various environments like the Internet of Things (IoT), Cyber-Physical Systems (CPSs), Software Defined Networks (SDNs), or Critical infrastructure Supervisory Control and Data Acquisition (SCADA).
Although not restricted to a particular environment, we strictly include only datasets containing network traces. Therefore, datasets without activity observable on a computer network (i.e., packets or flows) were excluded. As a result, the survey did not consider datasets that include only environment-specific features like sensor values typical for CPS or SCADA.
Given this scope, we emphasize that this survey is not supposed to be an exhaustive list of all available IDS datasets and, thus, should not serve as the sole data reference for specific use cases. Instead, we encourage consulting other surveys focused on specific cybersecurity subdomains as needed.
@@ 3 Domain-Specific Properties and Their Impact on Data
Despite its rapid adoption in research, ML penetrated into practical cybersecurity solutions much slower than in other domains, such as computer vision or natural language processing, which have been actively using ML for over a decade [17]. In fact, only a fraction of real-world network monitoring and threat detection teams relied on machine learning in early 2020s [18].
The discrepancy between academic research and the actual deployments of ML-based NIDSs was discussed by Sommer and Paxson [13], who noticed that operators were reluctant to adopt ML-based anomaly detectors despite high research efforts within the area. This fact was attributed to fundamental differences between NID and other ML-application domains, driven by factors like very high cost of errors, enormous data variability, semantic gap (ML explainability issues), and difficulties with evaluation – including the lack of benchmark data. Although published more than 10 years ago, most of these issues arose from the specific properties of cybersecurity and computer networking domains, affecting the research and development efforts to this day. This section elaborates on these problems and their impact on overall NID data availability and quality.
@@ 3.1 When Hell Comes to Earth: Unpleasant Properties of the Network Intrusion Detection Domain
    Network traffic diversity and dynamicity (intra-network variability): Computer networks are highly dynamic by nature. Network environments constantly change as new devices are connected to the network, new applications and protocols are developed, or even new networking paradigms, such as the Internet of Things (IoT), are introduced. The temporal aspect further introduces periodicity patterns based on day-and-night cycles and office hours [20].
@@
    Adversarial environment: In addition to natural networks’ dynamicity, the NID domain has to assume the implicit presence of adversaries. Due to the endless cybersecurity arms race, attacks continuously evolve, and new vulnerabilities are dicovered. Moreover, the attackers may also attempt to bypass detection via adversarial samples crafted with obfuscation techniques or tiny perturbations of the input data [21].
@@
    In NID, even domain experts might struggle to verify ground truth, whereas attack detectors might return different and even contradicting results [26]. These facts also disable crowdsourcing annotation, a common way to obtain large corpora of labeled data [27]. Finally, even all the traffic of infected hosts cannot be considered malicious, as some might correspond to legitimate activities (e.g., ARP messages) [23]. For these reasons, accurately labeling network data is difficult and often infeasible.
@@
    Data confidentiality: Captures from real networks might contain highly sensitive and confidential information, such as IP addresses or packet payloads, significantly limiting their sharing. Although anonymization techniques [28] like IP address remapping or payload stripping can be used, they decrease the data’s value – e.g., payload analyzers NIDSs cannot be used.
Given the challenges of labeling and data sharing, the scarcity of high-quality real-world data becomes understandable. However, the problem persists even when such data are collected and labeled. Since data is captured at a specific point in time, future changes in the underlying distribution (data drift) or a relationship between input features and the output labels (concept drift) violate the i. i. d. assumption. This phenomenon, known as model drift, leads to a gradual decline in a prediction model’s performance over time. Since NID is adversarial and highly dynamic, model drifting, amplified by the low tolerance of errors, is considered a major obstacle in ML utilization in real-world scenarios. While approaches like incremental [29], lifelong [30], or active learning [31] have been proposed to mitigate drift by leveraging newly observed data for model updates, applying ML for real-world NIDSs remains heavily constrained.
@@ 3.2 Limitations of Existing Datasets
    Timeliness: The direct consequence of constantly evolving network and threat landscapes is data timeliness. Generally, it cannot be assumed that the behavior collected during a certain period will remain immutable over time [32]. Therefore, static datasets become less representative of real-world scenarios as they age [33], affecting the soundness of the evaluation results. Although continual updates with the newest traffic patterns would mitigate this issue, immutable (static) datasets are required for fair benchmark comparisons between systems.
@@
    Testbed limitations: Due to the limitations of real-world data, most public datasets contain emulated traffic captured in a lab testbed environment. Despite some benefits of such an approach, its data are not guaranteed to resemble realistic network behavior and threats [36]. Such datasets tend to model over-simplified network conditions, failing to capture their real-world complexity and diversity [33, 8]. The lack of realism in benchmark data is considered a major obstacle in deploying anomaly-based NIDSs in real-world production environments [13, 32]. For instance, Catillo et al. [37] have shown that many attacks from DoS public datasets do not resemble real-world characteristics due to being inefficient (not achieving their malicious intent) or assuming non-realistic environment configurations.
@@
    Class imbalance: Modern machine learning classifiers are trained to maximize predictive accuracy by minimizing misclassification errors [38]. Such a setting usually assumes roughly balanced classes, i.e., with similar prior probabilities. However, cyberthreat detection domains like NID typically deal with a huge imbalance in favor of the negative (benign) class, as only a few attack instances are observed along a vast amount of non-malicious communication [33]. Although testing data should resemble real-world distribution to prevent spatial bias [39], training a supervised classifier should be done on approximately balanced data. Nevertheless, most datasets are imbalanced, leaving to address the issue by their users [40].
@@
    Documentation: The dataset’s accompanying metadata and documentation are crucial in determining the data value and its suitability for particular use cases. However, datasets often lack the documentation completely or omit necessary details. For this reason, the soundness of conclusions drawn by the data users is significantly hampered.
@@
    Lack of explicit train-test splits: Many datasets are not explicitly split for training and testing. As a result, their users create arbitrary splits, causing their methods to be incomparable, as each may be trained and evaluated on a different data subset. Furthermore, applying methods like cross-validation or partitioning the data without respecting time dependencies may lead to temporal bias and inflated evaluation performance [15, 39].
@@
    Feature set discrepancies: As shown in our survey, many dataset authors propose their own feature sets, typically based on network flows. While some provide only 10 – 15 essential features, others offer over 100 advanced statistics. Several newer datasets adopt the CICFlowMeter tool [41] for feature extraction. An attempt to unify feature sets was made by Sarhan et al. [42, 43], who proposed extracting 43 NetFlow v9 features using nProbe [44] flow collector. Despite these efforts, the community has not yet agreed on a unified, widely accepted feature set, hampering the transferability and generalizability between datasets with different features.
@@
    Limited scope: Due to the diversity of network technologies and cyberthreats, assembling a comprehensive dataset that covers the entire NID landscape is impractical. While specialized datasets (e.g., IoT) have been released recently, many scenarios and intrusions remain underrepresented. In particular, Hindy et al. [46] note that public datasets cover only 33% of known attacks. Given that new attacks and vulnerabilities continually emerge, a NIDS benchmarked on a static dataset will never be able to fully comprehend all potential threats in the wild.
@@
    Incorrect labeling: Due to networking dynamicity and the common methods for traffic labeling (based on the attack time and the source and destination addresses of the attacker/victim [47]), entries in NIDS datasets are prone to mislabeling. The issue has been demonstrated by Liu et al. [48] and Flood et al. [8] through analyses of several popular benchmark datasets. Such mislabeling can result in unrealistic results and benchmark destabilization [49].
  11. 11)
    Methodically wrong data handling: In addition to mislabeling, several other factors, primarily originating from human error in data collection and processing, affect the final dataset. A common issue we encountered relates to improper timestamping, as some datasets contain incorrect timestamps or time inconsistencies of events (e.g., attacks) with the documentation. We also observed disordered packets, undocumented capture gaps, and capture interruptions, resulting in improperly terminated files. Data processing issues, often due to bugs in feature extractor software [33], include missing values, duplicated features or entries, and zero-variance features (e.g., all zeros). These issues are frequently unreported, creating ambiguity for users and undermining the soundness of their conclusions.
While many discussed limitations arising from the domain-specific properties are challenging to address, others (documentation, lack of explicit splits, feature set discrepancies, reproducibility, and data handling errors) are entirely within the control of dataset authors. Even when a limitation cannot be fully mitigated, proper documentation reduces ambiguity and significantly enhances data usability. For this reason, we consider documentation a crucial pillar of every dataset. However, most do not discuss any limitations at all. Supposing that authors do not release inherently flawed records on purpose, we theorize that they often neglect thorough analysis of their collected data.
From the perspective of data users, understanding these limitations aids in efficient data selection, proper handling, and avoiding evaluation errors. In Section 6, we discuss recommendations to minimize or entirely prevent their impact.
@@ 4 Network Intrusion Data and Its Properties
Depending on the dataset goal and capture attributes, we characterize network intrusion data by various properties that influence their use cases and the soundness of methods utilizing them. In this survey, we collected 13 essential properties to provide an unbiased, comprehensive view and comparison of available datasets for NID purposes. A portion of them was adopted from Ring et al.’s survey [14].
We split the extracted information into five categories: General information, Nature of data, Data volume, Network properties, and Evaluation. This section elaborates on these categories and individual properties to aid in data understanding, selection, as well as facilitate the evaluation of datasets not included in this survey in future research. Their summary is provided in Table 1.
Table 1: Summary of 13 collected properties of the surveyed NID datasets split into five categories. The table briefly describes each property, and the fourth column (“Possible values”) is a comma-separated list of values a particular property can attain. Curly brackets symbolize a set from which a single value can be drawn, whereas $\times$ is a Cartesian product of such sets in cases when multiple elements define a single property. Ampersand (“&”) symbol is occasionally used to combine multiple values within a single entry.
@@ 4.1 General Information
When selecting a dataset, our primary focus is on the scenarios it includes, along with other details like the data collection period, the types of present attacks, and the environment used to perform the capture:
Year: The specific year(s) or period of data collection. Since many datasets are published months or years after their collection, reporting the actual collection date is crucial for understanding the captured threat landscape and traffic characteristics (e.g., popular applications). For continuously ongoing captures, such as MAWILab [50], we denote the value as Now.
Focus: This property summarizes the dataset scenario and its potential use cases. Based on the focus, we distinguish between general- or special-purpose datasets. General-purpose datasets (General) are captured in traditional local area or enterprise networks with various intrusive scenarios. Depending on the nature of the data, we include IDS (both host and network information) or NIDS (only network data are present) identifiers. In contrast, special-purpose datasets focus on a particular attack like DDoS or a specific environment such as IoT.
Normal: Indicates whether normal (benign) traffic is present and how it was generated. If absent, the value is recorded as no. Otherwise, we distinguish between data obtained from real, emulated (emul), or synthetic (synth) environments. Since various definitions of these terms are found in the literature (especially, emulated and synthetic terms are often used interchangeably), we establish the following definitions for this survey:
  • real – Traffic originating from real-world operational networks captured using tools like tcpdump or nfdump. It reflects genuine network interactions with purposes beyond traffic generation. Since it contains real information, it must be anonymized. Additionally, real data cannot be declared benign with absolute confidence, as undetected attacks might be present. Instead, it is referred to as background traffic without evident anomalous or malicious behavior.
@@
In contrast, emulated and synthetic traffic offer greater flexibility, lower collection costs, and more reliable labeling. However, as outlined in Section 3.2, their crucial limitation is realism. Due to the complexity of network traffic, replicating real-world scenarios in a laboratory environment is very challenging [33]. Although several works have tried to tackle this issue, e.g., $\beta$-profile system [54], realistic simulation of network behavior remains an important research direction [46].
@@
The attack origin environment is defined the same as for normal traffic – real, emul-{p, v, pv}, and synth. Real attacks correspond to traffic with actual malicious intent captured in real-world production networks, i.e., by detecting and sampling an ongoing attack, as in the Booters dataset [57]. Emulated attacks are launched in controlled environments just to simulate attacks’ behavior, usually using the same tools as attackers (e.g., in Kali Linux [58]). Finally, synthetic attacks were not captured in a network but rather generated by specialized algorithms.
@@
The environment origin is followed by a value in brackets indicating the number of distinguishable attacks. This represents the number of malicious classes clearly separable in the data (either on a feature-based level based on labels or via raw packet filtering). For datasets like CUPID [60] that declare multiple attacks but provide only binary labels, i.e., distinguishable anomalies, the value is listed as one. Datasets with unspecified attack counts, e.g., MAWILab [50], have no associated value.
In the third part, we summarize the attack types to clarify the dataset’s objectives and associated traffic. Although existing cyberattack taxonomies [61, 62] could be used, we developed a custom scheme (Table 2) to provide a concise summary of common attack events in NID datasets. Therefore, rather than a comprehensive taxonomy, it serves as a quick reference for identifying attacks within datasets. In contrast to other taxonomies that aim for completeness and specificity [61], our scheme intentionally includes the Unknown category to account for potentially undiscovered attacks, as well as the Other category for less common ones. Each attack type is then represented by an acronym (e.g., “D” for a (D)DoS attack), with multiple types forming a single acronym string.
@@
Table 2: The scheme for a quick reference of attacks present in the reviewed datasets. Each attack category is assigned an acronym, such as “B" for Brute-force. A sequence of such acronyms (e.g., BCO) then represents the presence of given types of attacks within the dataset.
@@
C&C Botnet traffic C Command and Control (C&C) traffic is used by infected machines (bots) to communicate and receive orders from the master server within the client-server botnet architectures.
@@
Vulnerability exploitation V The act of exploiting various vulnerabilities, such as SQL injection or buffer overflow, typically to gain remote access or escalate privileges on the target machine.